# Experimentation on Endogenous Graphs

### v1

@@ Abstract
We study experimentation under endogenous network interference. Interference patterns are mediated by an endogenous graph, where edges can be formed or eliminated as a result of treatment. We show that conventional estimators are biased in these circumstances, and present a class of unbiased, consistent and asymptotically normal estimators of total treatment effects in the presence of such interference. Our results apply both to bipartite experimentation, in which the units of analysis and measurement differ, and the standard network experimentation case, in which they are the same.
Keywords. Experimentation, interference, networks.
@@ 1.1 Background
In many settings, the fundamental no-interference assumption common in experimental analysis (Cox, 1958; Rubin, 1980) is implausible, and there are strong grounds for believing that the treatment assigned to one unit may affect outcomes of other units. This is especially true when outcomes are in part determined by peoples’ interactions with each other, as is often the case in public health, education, voting, social media, and other social domains (Calvó-Armengol et al., 2009; DiMaggio and Garip, 2012; Halloran and Hudgens, 2016; Taylor and Eckles, 2018). A large literature has arisen developing the theory of estimation and inference for various causal quantities of interest under these conditions.
Typically, theoretical and empirical analysis proceeds assuming that the interference structure is known—that is, the researcher has prior knowledge of the set of units which might be affected by the treatment of any given unit. This is usually expressed in terms of a known “exposure mapping” or “effective treatment” function, which determines the level of treatment exposure a unit receives as a function of the treatment assignment vector to all units, or in terms of a “partial interference” or “neighborhood interference” assumption, according to which which interference takes place only within groups, where the groups are known ex-ante (Hudgens and Halloran, 2008; Tchetgen and VanderWeele, 2012; Manski, 2013; Liu and Hudgens, 2014; Aronow and Samii, 2017; Forastiere et al., 2022). This is a natural starting point, and it substantially facilitates the development of treatment effect estimators and the characterization of their behavior.
In practice, however, this interference structure is rarely known (Egami, 2021; Sävje, 2024). Analysts may observe past user interactions, and construct interference graphs from these data (Aral et al., 2009; Bakshy et al., 2012; Bond et al., 2012; Coppock et al., 2016; Harshaw et al., 2021; Karrer et al., 2021). However these graphs may be only crude approximations to the underlying interference graph of interest. Complicating matters further, those data may themselves be causally affected by the treatment of interest. In online social networks, one might use friendship or connection graphs as proxies capturing which other users’ treatments affect a given user. Similarly, in the case of bipartite experimentation in a marketplace, one might use signals of past engagement to construct the bipartite interference graph connecting buyers and sellers. In contrast to the true, unobserved and non-stochastic interference graph, such convenience graphs can be affected by the particular realization of treatment assignments to the randomization units. Even if such graphs do capture some valuable information about how treatments affect outcomes, treating them as capturing the true interference structure may be problematic. The analyst is presented with a uncomfortable choice—either use a graph constructed only on the basis of pre-treatment data, which may omit relevant information about how treatments affect outcomes, or incorporate post-treatment data too, which can lead to complex biases in treatment effect estimators.
@@ 1.2 Our Contributions
Motivated by this gap between theory and practice, this paper takes a different approach, explicitly allowing for the observed graph available to the analyst to be determined by treatment assignment. We work in a general bipartite setting, in which units of analysis and randomization are allowed to differ. We propose the notion of an endogenous bipartite graph which may be treatment-dependent, and which contains information on interference from randomization to analysis units. Generalizing the outcome model of Harshaw et al., (2021) to this setting, we give an unbiased, consistent and asymptotically normal estimator of the total treatment effect, as well as a consistent testing procedure for the sharp null of no treatment effect. Our identification and estimation approach relies on an treatment-invariant “anchor” subgraph, from which instrumental variable estimates of unit-level treatment responses can be obtained.
Our results apply directly to bipartite experiments, which are of growing interest in their own right. Units of analysis and randomization may differ because it is impractical or impossible to measure outcomes of interest at the level at which treatments are being assigned, as in Zigler and Papadogeorgou, (2021), who study how interventions to reduce air pollution from power plants affect cardiovascular hospitalizations. Alternatively, there may be no conceptual difficulty in defining outcomes at the level of the randomization unit, but accounting for the bipartite structure is a natural way to allow for cross-unit interference. Harshaw et al., (2021) give the example of two-sided marketplaces: comparing sales of a products which do and do not receive random discounts may mislead regarding the global effect of discounting all goods, as buyers’ demand for one good depends on the prices of other goods. Causal inference problems with this structure are common in social networks, recommender systems, digital advertising, and multi-sided platforms broadly (Chawla et al., (2016); Gilotte et al., (2018); Pouget-Abadie et al., (2019); Nandy et al., (2021); Johari et al., (2022); Bajari et al., (2023); Shi et al., (2024)).
We show how our bipartite results apply with minimal modifications to the standard case of unipartite interference, in which the units of analysis and randomization are the same. In this setting it is natural to separately estimate the direct effect of a unit being treated, and the indirect effect coming from the treatment of neighboring units. We combine unbiased inverse-propensity weighted estimators of the direct effect, while estimating the indirect effect using the previously developed logic for bipartite estimators. Unbiasedness, consistency and asymptotic normality follow as corollaries to our results for bipartite graphs.
@@ 1.3 Related Work
Motivated by the observation that applying standard methods often requires experimenters to make questionable assumptions on the interference structure, an emerging literature explores relaxing this requirement, and allowing for misspecification of the interference structure (Aronow and Samii, 2017; Eckles et al., 2017; Wang et al., 2020; Leung, 2022; Sävje, 2024). Another closely related literature studies causal inference on unknown graphs (Basse and Airoldi, 2018; Chin, 2018; Egami, 2021; Sävje et al., 2021; Cortez et al., 2022; Yu et al., 2022; Shirani and Bayati, 2023). Our work shares the same concern that being able to correctly specify the interference structure is implausible. In contrast to the existing literature, our focus is on the case of endogenous misspecification, where treatments do indeed propagate to outcomes through the observed edges, but the set of observed edges is allowed to depend on the treatment assignment realization. To the best of our knowledge, these are the first available results on treatment effect estimation on endogenous graphs of any kind, whether bipartite or unipartite. Several important questions remain open, including consistent variance estimation for the total treatment effect estimator.
@@ 1.4 Outline of the Paper
The rest of the paper is organized as follows. Section 2 introduces the basic setup for endogenous bipartite interference graphs and discusses the edge endogeneity bias. Section 3 proposes several models for endogenous edge formation models and statistical tests to detect them. Section 4 includes our main estimator, and for which we demonstrates unbiasedness, consistency and asymptotic normality under certain assumptions. We also propose a statistical test for the sharp null with asymptotic power equal to $1$. Section 5 extends the results to unipartite graphs, and Section 6 concludes.
@@ 2 Setup
There are $n_{a}$ analysis units $a\in\mathcal{A}$ and $n_{r}$ randomization units $r\in\mathcal{R}$. We denote the treatment assignment random vector by $\mathbf{T}=(T_{1},T_{2},\ldots,T_{n_{r}})$. For each analysis unit $a$, the potential outcome function $Y_{a}:\{0,1\}^{n_{r}}\to\mathbb{R}$ maps treatment assignment vectors to outcomes. Throughout, we consider the object of interest to be the total treatment effect (TTE), $\mu=\frac{1}{n_{a}}\sum_{a}[Y_{a}(\mathbf{1})-\sum_{a}Y_{a}(\mathbf{0})]$. All randomness comes from treatment assignment. In the usual bipartite experimentation setting, the graph is defined so that an edge exists between an analysis unit and a randomization unit if the potential outcomes of the analysis unit depend on the treatment of the randomization unit. The following definition describes this more formally.
@@
In this definition the adjacency matrix $E$ restricts how each analysis unit’s outcomes can vary in response to treatment assignments, and furthermore is invariant to $\mathbf{T}$. Beyond the trivial example of a fully connected graph, it is rarely plausible to assume that this knowledge is available in applications. We take a different approach, motivated by practical settings in which the analyst observes and wishes to make use of a bipartite graph which contains some important information about how treatments affect outcomes, but cannot plausibly satisfy the definition above. In particular, we will allow for the bipartite graph to be determined by the treatment assignment.
Formally, potential outcomes for edges are a function of treatment assignments. Each pair $(a,r)\in\mathcal{A}\times\mathcal{R}$ has a corresponding edge potential outcome function $E_{ar}:\{0,1\}^{n_{r}}\to\{0,1\}$, i.e. specifying whether the corresponding edge exists as a function of the treatment assignment vector. Let $\mathcal{R}_{a}(\mathbf{T})=\{r\in\mathcal{R}:E_{ar}(\mathbf{T})=1\}$ denote the randomization units $a$ is connected to under the treatment $\mathbf{T}$. We assume that as long as a treatment assignment does not change the randomization units that analysis unit $a$ is connected to, and does not change the treatment assignments to those randomization units, it cannot change $a$’s outcome. This gives the definition of an endogenous bipartite interference graph. Let $\mathcal{R}_{a}(\mathbf{T})=\{r\in\mathcal{R}:E_{ar}(\mathbf{T})=1\}$ denote the randomization units that analysis unit $a$ is connected to under treatment assignment $\mathbf{T}$.
@@
We assume throughout that for all $(a,r)\in\mathcal{A}\times\mathcal{R}$ the functions $E_{ar}(\cdot)$ are unknown, but their realizations given the treatment assignment $E_{ar}(\mathbf{T})$ are observed. Some observations on the relation between bipartite vs. unipartite and endogenous vs. exogenous interference graphs follow.
@@
Definition 2 also encompasses the case of endogenous unipartite experimentation, which is the special case of $\mathcal{A}=\mathcal{R}$. In this case we typically assume $E_{aa}=1$ for all $a$, so that each unit’s treatment assignment is allowed to affect itself. The interference graph is allowed to be directed, in the sense that $E_{aa^{\prime}}$ need not necessarily equal $E_{a^{\prime}a}$ for $a\neq a^{\prime}$.
@@
Given any endogenous bipartite interference graph $(\mathcal{A},\mathcal{R},E)$, it is always possible to find some (exogenous) bipartite interference graph $(\mathcal{A},\mathcal{R},E^{\prime})$ consistent with $(\mathcal{A},\mathcal{R},E)$. Trivially, we can take $E^{\prime}$ to be the fully connected bipartite graph $\mathcal{A}\times\mathcal{R}$. Somewhat less trivially, we can define $E^{\prime}$ as the set of edges $(a,r)$ such that there exists a treatment realization $\mathbf{T}$ satisfying $E_{ar}(\mathbf{T})=1$. In principle conventional methods for causal inference on networks could be applied on the graph $(\mathcal{A},\mathcal{R},E^{\prime})$. This observation is of limited value to practitioners, if $E^{\prime}$ is unobserved and they only observe a particular, treatment-dependent realization $E(\mathbf{T})$ of endogenous edges.
It is unreasonable to expect graphs constructed from behavioral signals to satisfy Definition 1 if they are likely to be causally affected by treatments. But they may be consistent with Definition 2. For example, for a user-item graph in an online marketplace, we may define an edge as corresponding to the item being shown to the user, which is information that can be logged. While this graph may be causally affected by item-level treatments, we may still be prepared to assume that a treatment which does not affect the items surfaced to a user, or how those items themselves are treated, cannot change that user’s outcomes. Similarly, a experiment on Facebook groups may affect which users are members of which groups, but it may be plausible that a treatment assignment to Facebook groups will not change a user’s behavior, as long as it does not change the Facebook groups a user is a member of, or the treatment assignments to those groups.
@@
Treating the realized graph $E(\mathbf{T})$ observed in the setting of Definition 2 as fixed can bias otherwise unbiased estimators, including the Horvitz–Thompson inverse propensity weighted estimator (Horvitz and Thompson, (1952)). We assume the treatment assignments are independent Bernoulli random variables with probability $p$. The Horvitz–Thompson estimator of the TTE is
$\hat{\mu}_{\text{HT}}$ $=\frac{1}{n_{a}}\sum_{a=1}y_{a}\left[\frac{\prod_{r\in\mathcal{R}_{a}(\mathbf{T})}T_{r}}{p^{|\mathcal{R}_{a}(\mathbf{T})|}}-\frac{\prod_{r\in\mathcal{R}_{a}(\mathbf{T})}(1-T_{r})}{(1-p)^{|\mathcal{R}_{a}(\mathbf{T})|}}\right],$
@@
where the term in the square brackets is defined to be zero if $\mathcal{R}_{a}(\mathbf{T})=\emptyset$. The first term simply sums up outcomes $y_{a}$ for the “fully treated” analysis units, i.e. those units $a$ such that all their corresponding randomization units are treated ($\prod_{r\in\mathcal{R}_{a}(\mathbf{T})}T_{r}=1$), and rescales by the factor $p^{-|\mathcal{R}_{a}(\mathbf{T})|}$ to account for the probability of that event. This gives the estimate of the average outcomes under $\mathbf{T}=\mathbf{1}$, and analogously the second term gives the estimate of average outcomes under $\mathbf{T}=\mathbf{0}$. By standard arguments following Aronow and Samii, (2017), this is an unbiased estimate of the TTE under graph exogeneity, so that $\mathcal{R}_{a}(\mathbf{T})$ is constant with respect to $\mathbf{T}$.
@@
Under graph endogeneity, this unbiasedness property no longer holds. We illustrate mechanisms by which bias can arise in the following minimal examples. There is a single randomization unit in each, implying the Horvitz–Thompson estimator and the exposure reweighted linear (ERL) estimator of Harshaw et al., (2021) are identical, and so these examples also demonstrate bias in the ERL estimator.
@@
Example 1. =$\mathcal{R}=\{r_{1}\}$ and $\mathcal{A}=\{a_{1}\}$. Unit $r_{1}$ forms an edge with $a_{1}$ if and only if $T_{1}=1$. Outcomes for $a_{1}$ are $y\neq 0$ regardless of the treatment assignment, and $T_{1}\sim\text{Bernoulli}(1/2)$. The Horvitz–Thompson estimator is $\frac{T_{1}y}{1/2}$. This is $y$ in expectation, which is biased for the TTE of zero. In this example the number of edges varies with the treatment assignment, but that is not necessary for bias to exist, as the next example shows.
@@
Example 2. =$\mathcal{R}=\{r_{1}\}$ and $\mathcal{A}=\{a_{1},a_{2}\}$. Unit $r_{1}$ forms an edge with $a_{1}$ if $T_{1}=1$, and with $a_{2}$ otherwise. Outcomes for $a_{1}$ and $a_{2}$ are $y_{1}$ and $y_{2}$ regardless of the treatment assignment, and $T_{1}\sim\text{Bernoulli}(1/2)$. The Horvitz–Thompson estimator is $\frac{T_{1}y_{1}}{1/2}-\frac{(1-T_{1})y_{2}}{1/2}$ which has expectation $y_{1}-y_{0}$, and again is biased for the TTE of zero. In this example, edge formation induces correlation between treatment intensity and analysis unit potential outcomes, generating bias.
@@
The preceding examples distinguish between randomization and analysis units and so relate to bipartite experimentation. Graph endogeneity can also bias treatment effect estimators in the familiar case of unipartite network interference where this distinction does not exist, as the next example shows.
@@
Example 3. $\mathcal{R}=\mathcal{A}=\{a_{1},a_{2}\}$. Outcomes for $a_{1}$ and $a_{2}$ are $y_{1}$ and $y_{2}$ regardless of the treatment assignment, and $T_{1}$, $T_{2}$ are independent $\text{Bernoulli}(1/2)$ draws. A directed edge runs from $a_{1}$ to $a_{2}$ if and only if $T_{1}=1$. The Horvitz-Thompson estimator is
$\hat{\mu}_{\text{HT}}=\begin{cases}-y_{1}/2-y_{2}/2&\text{if $T_{1}=T_{2}=0$}\\ -y_{1}/2+y_{2}/2&\text{if $T_{1}=0,T_{2}=1$}\\ y_{1}/2&\text{if $T_{1}=1,T_{2}=0$}\\ y_{1}/2+y_{2}/4&\text{if $T_{1}=T_{2}=1$}\\ \end{cases}$
which has expectation $y_{2}/16$, and is biased for the TTE of zero.

### v2

@@ Abstract
We study experimentation under endogenous network interference. Interference patterns are mediated by an endogenous graph, where edges can be formed or eliminated as a result of treatment. We show that conventional estimators are biased in these circumstances, and present a class of unbiased, consistent and asymptotically normal estimators of total treatment effects in the presence of such interference. We show via simulation that our estimator outperforms existing estimators in the literature. Our results apply both to bipartite experimentation, in which the units of analysis and measurement differ, and the standard network experimentation case, in which they are the same.
@@ 1.1 Background
In many settings, the fundamental no-interference assumption common in experimental analysis (Cox 1958, Rubin 1980) is implausible, and there are strong grounds for believing that the treatment assigned to one unit may affect outcomes of other units. This is especially true when outcomes are in part determined by peoples’ interactions with each other, as is often the case in public health, education, voting, social media, and other social domains (Calvó-Armengol et al. 2009, DiMaggio and Garip 2012, Halloran and Hudgens 2016, Taylor and Eckles 2018). A large literature has arisen developing the theory of estimation and inference for various causal quantities of interest under these conditions.
Typically, theoretical and empirical analysis proceeds assuming that the interference structure is known—that is, the researcher has prior knowledge of the set of units which might be affected by the treatment of any given unit. This is usually expressed in terms of a known “exposure mapping” or “effective treatment” function, which determines the level of treatment exposure a unit receives as a function of the treatment assignment vector to all units, or in terms of a “partial interference” or “neighborhood interference” assumption, according to which interference takes place only within groups, where the groups are known ex-ante (Hudgens and Halloran 2008, Tchetgen and VanderWeele 2012, Manski 2013, Liu and Hudgens 2014, Aronow and Samii 2017, Forastiere et al. 2022). This is a natural starting point that greatly facilitates the development of treatment effect estimators and the characterization of their behavior.
In practice, the interference structure is rarely known (Egami 2021, Sävje 2024). Analysts often rely on past user interactions to construct proxy interference graphs (Aral et al. 2009, Bakshy et al. 2012, Bond et al. 2012, Coppock et al. 2016, Harshaw et al. 2021, Karrer et al. 2021). These graphs—such as friendship networks in social media or engagement-based links in marketplaces—are typically crude approximations of the true, unobserved interference graph. [^1] For example, social network users frequently interact with “unconnected nodes” by being recommended to join new groups or follow new users. Complicating matters further, they may themselves be causally affected by the treatment of interest. The analyst is presented with an uncomfortable choice—either use a graph constructed only on the basis of pre-treatment data, which may omit relevant information about how treatments affect outcomes, or incorporate post-treatment data too, which can lead to complex biases in treatment effect estimators.
@@ 1.2 Our Contributions
Motivated by this gap between theory and practice, this paper takes a different approach, explicitly allowing for the observed graph to be determined by treatment assignment. We work in a general bipartite setting, in which units of analysis and randomization are allowed to differ. We propose the notion of an endogenous bipartite graph which may be treatment-dependent, and which contains information on interference from randomization to analysis units. Generalizing the outcome model of Harshaw et al. (2021) to this setting, we give an unbiased, consistent and asymptotically normal estimator of the total treatment effect, as well as a consistent testing procedure for the sharp null of no treatment effect. Our identification and estimation approach relies on a treatment-invariant “anchor” subgraph, from which instrumental variable estimates of unit-level treatment responses can be obtained.
Our results apply directly to bipartite experiments, which are of growing interest in their own right. Units of analysis and randomization may differ because it is not practical or impossible to measure outcomes of interest at the level at which treatments are being assigned (Zigler and Papadogeorgou 2021). Alternatively, there may be no conceptual difficulty in defining outcomes at the randomization unit-level, but accounting for the bipartite structure is a natural way to allow for cross-unit interference. Causal inference problems with this structure are common in social networks, recommender systems, digital advertising, and multi-sided platforms broadly (Chawla et al. (2016), Gilotte et al. (2018), Pouget-Abadie et al. (2019), Harshaw et al. (2021), Nandy et al. (2021), Johari et al. (2022), Bajari et al. (2023), Shi et al. (2024)).
Our framework extends naturally to the standard unipartite setting, where the units of analysis and randomization are the same. There, we combine unbiased inverse-propensity weighted estimators of the direct effect, while estimating the indirect effect using the previously developed logic for bipartite estimators. Unbiasedness, consistency and asymptotic normality follow as corollaries to our results for bipartite graphs.
@@ 1.3 Related Work
Two subareas of network science are especially related to this paper: models of network formation and estimation of peer effects (see An (2011), Chandrasekhar (2016), Bramoullé et al. (2020) for overviews). Particularly relevant is an emerging body of work that questions standard assumptions about interference structure and allows for misspecification of the interference graph (Aronow and Samii 2017, Eckles et al. 2017, Wang et al. 2020, Leung 2022, Sävje 2024).
Related work also tackles causal inference on unknown graphs (Basse and Airoldi 2018, Chin 2018, Egami 2021, Sävje et al. 2021, Cortez et al. 2022, Yu et al. 2022, Shirani and Bayati 2023, Halloran and Hudgens 2016). Like these papers, we are concerned with the implausibility of correctly specifying the interference structure. However, we focus on the case of endogenous misspecification: treatment effects propagate along observed edges, but those edges may themselves depend on treatment assignment realizations. Notable recent exceptions that address this setting include Comola and Prina (2021), Gao (2024), Ryu (2024). Unlike Comola and Prina (2021), we estimate the effect of treating everyone vs no one, the policy-relevant estimand. Ryu (2024) takes a non-instrumental variable-based identification approach in the unipartite setting, using conditional means for outcome and edge formation with parameters assumed to be shared across units. Most closely related is Gao (2024), which also develops an instrumental variable strategy based on a pre-treatment graph, proceeding from quite different modeling assumptions (a unipartite graph with undirected, unweighted edges, and an outcome model with parameters shared across units).
To the best of our knowledge, in addition to providing novel insights into identification and estimation of treatment effects on endogenous unipartite graphs, these are also the first available results on treatment effect estimation for endogenous bipartite graphs. Several important questions remain open, including consistent variance estimation for the total treatment effect estimator.
@@ 1.4 Outline of the Paper
The rest of the paper is organized as follows. Section 2 introduces the basic setup for endogenous bipartite interference graphs and discusses the edge endogeneity bias. Section 3 proposes several models of endogenous edge formation. Section 4 includes our main estimator, and for which we demonstrate unbiasedness, consistency and asymptotic normality under certain assumptions. We also propose a statistical test for the sharp null with asymptotic power equal to $1$. Our theory is supported by simulations in Section 5, where we examine the biases of our estimator and two exposure reweighted linear estimators (Harshaw et al. 2021), and demonstrate the confidence interval coverage properties of our estimator based on a variance proxy. Section 6 extends the results to unipartite graphs, and Section 7 concludes. All proofs and additional details are included in the Supplementary material.
@@ 2 Setup
There are $n_{a}$ analysis units $a\in\mathcal{A}$ and $n_{r}$ randomization units $r\in\mathcal{R}$. The treatment assignment random vector is denoted by $\mathbf{T}=(T_{1},T_{2},\ldots,T_{n_{r}})$. For each analysis unit $a$, the potential outcome function $Y_{a}:\{0,1\}^{n_{r}}\to\mathbb{R}$ maps treatment assignment vectors to outcomes. Throughout, the object of interest is the total treatment effect (TTE), $\mu=\frac{1}{n_{a}}\sum_{a}[Y_{a}(\mathbf{1})-Y_{a}(\mathbf{0})]$. All randomness comes from treatment assignment. In the usual bipartite experimentation setting, the graph is defined so that an edge exists between an analysis unit and a randomization unit if the potential outcomes of the analysis unit depend on the treatment of the randomization unit. The next definition formalizes this.
@@
In this definition, the adjacency matrix $E$ restricts how each analysis unit’s outcomes can vary in response to treatment assignments, and furthermore is invariant to $\mathbf{T}$. Beyond the trivial example of a fully connected graph, it is rarely plausible to assume that this knowledge is available in applications. We take a different approach, motivated by practical settings in which the analyst observes and wishes to make use of a bipartite graph which contains some important information about how treatments affect outcomes, but cannot plausibly satisfy the definition above. In particular, we allow for the bipartite graph to be determined by the treatment assignment.
Figure 1: Endogenous graph example. Note that after the treatment is applied, edges (A-1) and (B-4) disappear, while new edges (B-1) and (C-3) emerge.
Formally, potential outcomes for edges are a function of treatment assignments. Each pair $(a,r)\in\mathcal{A}\times\mathcal{R}$ has a corresponding edge potential outcome function $E_{ar}:\{0,1\}^{n_{r}}\to\{0,1\}$, i.e. specifying whether the corresponding edge exists as a function of the treatment assignment vector. Let $\mathcal{R}_{a}(\mathbf{T})=\{r\in\mathcal{R}:E_{ar}(\mathbf{T})=1\}$ denote the randomization units $a$ is connected to under the treatment $\mathbf{T}$. We assume that as long as a treatment assignment does not change the randomization units that analysis unit $a$ is connected to and does not change the treatment assignments to those randomization units, it cannot change $a$’s outcome. This gives the definition of an endogenous bipartite interference graph.
@@
We assume throughout that for all $(a,r)\in\mathcal{A}\times\mathcal{R}$ the functions $E_{ar}(\cdot)$ are unknown, but their realizations given the treatment assignment $E_{ar}(\mathbf{T})$ are observed (see Figure 1 for illustration). Some observations on the relation between bipartite vs. unipartite and endogenous vs. exogenous interference graphs follow.
@@
Definition 2 also encompasses the case of endogenous unipartite experimentation, which is the special case of $\mathcal{A}=\mathcal{R}$. In this case, we typically assume $E_{aa}=1$ for all $a$, so that each unit’s treatment assignment is allowed to affect itself. The interference graph is allowed to be directed, in the sense that $E_{aa^{\prime}}$ need not necessarily equal $E_{a^{\prime}a}$ for $a\neq a^{\prime}$.
@@
Given any endogenous bipartite interference graph $(\mathcal{A},\mathcal{R},E)$, it is always possible to find some (exogenous) bipartite interference graph $(\mathcal{A},\mathcal{R},E^{\prime})$ consistent with $(\mathcal{A},\mathcal{R},E)$. Trivially, we can take $E^{\prime}$ to be the fully connected bipartite graph $\mathcal{A}\times\mathcal{R}$. Somewhat less trivially, we can define $E^{\prime}$ as the set of edges $(a,r)$ such that there exists a treatment realization $\mathbf{T}$ satisfying $E_{ar}(\mathbf{T})=1$. In principle, conventional methods for causal inference on networks could be applied on the graph $(\mathcal{A},\mathcal{R},E^{\prime})$. This observation is of limited value to practitioners, if $E^{\prime}$ is unobserved and they only observe a particular, treatment-dependent realization $E(\mathbf{T})$ of endogenous edges.
Graphs built from behavioral signals may not satisfy Definition 1 if they are likely to be causally affected by treatments, but they may still satisfy Definition 2. For instance, a user-item graph in an online marketplace where edges represent shown items, may be causally affected by item-level treatments. Yet we might still assume that user outcomes remain unchanged if neither the set of shown items nor their treatments change. Similarly, in a Facebook Groups experiment, user outcomes may plausibly be unaffected if a user’s group memberships and the treatments of those groups remain the same.
@@
Treating the realized graph $E(\mathbf{T})$ observed in the setting of Definition 2 as fixed can bias otherwise unbiased estimators, including the Horvitz-Thompson (HT) inverse propensity weighted estimator (Horvitz and Thompson (1952)). We assume that treatment assignments are independent Bernoulli random variables with probability $p$. The HT estimator of the TTE is
$\hat{\mu}_{\text{HT}}$ $=\frac{1}{n_{a}}\sum_{a=1}^{n_{a}}y_{a}\left[\frac{\prod_{r\in\mathcal{R}_{a}(\mathbf{T})}T_{r}}{p^{|\mathcal{R}_{a}(\mathbf{T})|}}-\frac{\prod_{r\in\mathcal{R}_{a}(\mathbf{T})}(1-T_{r})}{(1-p)^{|\mathcal{R}_{a}(\mathbf{T})|}}\right],$
@@
where the term in the square brackets is defined to be zero if $\mathcal{R}_{a}(\mathbf{T})=\emptyset$. The first term simply sums up outcomes $y_{a}$ for the “fully treated” analysis units, i.e. those units $a$ such that all their corresponding randomization units are treated ($\prod_{r\in\mathcal{R}_{a}(\mathbf{T})}T_{r}=1$), and rescales by the factor $p^{-|\mathcal{R}_{a}(\mathbf{T})|}$ to account for the probability of that event. This gives the estimate of the average outcomes under $\mathbf{T}=\mathbf{1}$, and analogously the second term gives the estimate of average outcomes under $\mathbf{T}=\mathbf{0}$. By standard arguments Aronow and Samii (2017), this is an unbiased estimate of the TTE under graph exogeneity, where $\mathcal{R}_{a}(\mathbf{T})$ does not vary with $\mathbf{T}$.
@@
Under graph endogeneity, this unbiasedness property no longer holds. We illustrate mechanisms by which bias can arise in the following minimal examples. There is a single randomization unit in each, implying the HT estimator and the exposure reweighted linear (ERL) estimator of Harshaw et al. (2021) are identical, so these examples also demonstrate bias in the ERL estimator.
Example 1.
@@
$\mathcal{R}=\{r_{1}\}$ and $\mathcal{A}=\{a_{1}\}$. Unit $r_{1}$ forms an edge with $a_{1}$ if and only if $T_{1}=1$. Outcomes for $a_{1}$ are $y\neq 0$ regardless of the treatment assignment, and $T_{1}\sim\text{Bernoulli}(1/2)$. The HT estimator is $\frac{T_{1}y}{1/2}$. This is $y$ in expectation, which is biased for the TTE of zero. In this example the number of edges varies with the treatment assignment, but that is not necessary for bias to exist, as the next example shows.
Example 2.
@@
$\mathcal{R}=\{r_{1}\}$ and $\mathcal{A}=\{a_{1},a_{2}\}$. Unit $r_{1}$ forms an edge with $a_{1}$ if $T_{1}=1$, and with $a_{2}$ otherwise. Outcomes for $a_{1}$ and $a_{2}$ are $y_{1}$ and $y_{2}$ regardless of the treatment assignment, and $T_{1}\sim\text{Bernoulli}(1/2)$. The HT estimator is $\frac{T_{1}y_{1}}{1/2}-\frac{(1-T_{1})y_{2}}{1/2}$ which has expectation $y_{1}-y_{2}$, and again is biased for the TTE of zero. In this example, edge formation induces correlation between treatment intensity and analysis unit potential outcomes, generating bias.
@@
The preceding examples distinguish between randomization and analysis units and so relate to bipartite experimentation. Graph endogeneity can also bias treatment effect estimators in the case of unipartite network interference where this distinction does not exist, as the next example shows.
Example 3.
@@
$\mathcal{R}=\mathcal{A}=\{a_{1},a_{2}\}$. Outcomes for $a_{1}$ and $a_{2}$ are $y_{1}$ and $y_{2}$ regardless of treatment assignment, and $T_{1}$, $T_{2}$ are independent $\text{Bernoulli}(1/2)$ draws. A directed edge runs from $a_{1}$ to $a_{2}$ if and only if $T_{1}=1$. The HT estimator is
[…]
with expectation $y_{2}/16$, and is biased for the TTE of zero.